Please help, I don't even know what kind of question this is. 6/x = 8/x - 1/11
If I knew what it was I might could figure it out. Any hints?

Answer by edjones(8007)

6/x = 8/x - 1/11
66=88-x Multiply each side by 11x
x=22
.
Ed